(19) Assume Xi and X2 are independent random variables with a common proba- bility density function: f(x) 0<x<5 , Determine the cumulative distribution fun G(s), where Y = max(X1, X2) 1 5 The circled answer is wrong Just show steps to arrive at a correct answer please.
Answer #1

here cumulative distribution function:G(y)= P(Y<y) =P(max(X1,X2)<y)

=P(X1<y)*P(X2<y)=(y/5)*(y/5) =y2/25

option B is correct